OpenStudy (anonymous):

Which value of x makes the equation true? x – 18.7 = –12 x =

OpenStudy (anonymous):

@MikuxLuka

OpenStudy (shreenithaa):

x-18.7=-12 +18.7= +18.7 x=6.7

OpenStudy (anonymous):

so 6.7?

OpenStudy (shreenithaa):

yup

OpenStudy (anonymous):

thanks

OpenStudy (shreenithaa):

ur welcome